# Efficient estimation of error bounds for quantum multiparametric imaging with constraints

### v2

@@ Abstract
Advanced super-resolution imaging techniques require specific approaches for accurate and consistent estimation of the achievable spatial resolution. Fisher information supplied to Cramér-Rao bound (CRB) has proved to be a powerful and efficient tool for resolution analysis and optical setups optimization. However, the standard CRB is not applicable to constrained problems violating the unbiasedness condition, while such models are frequently encountered in quantum imaging of complex objects. Complimentary to the existing approaches based on modifying CRB, we propose a practical algorithm for approximate construction of a modified Fisher information matrix, which takes the constraints into account and can be supplied to the standard CRB. We demonstrate the efficiency of the proposed technique by applying it to 1-, 2-, and multi-parameter model problems in quantum imaging. The approach provides quantitative explanation of previous results with successful experimental reconstruction of objects with the spatial scale smaller than the theoretical limit predicted by the standard CRB.
@@ I Introduction
Efficiency of CRB for resolution analysis on the one hand and typicality of constrained problems in practical imaging on the other hand make it tempting to reconcile them. One of the available approaches is based on constructing the modified CRB [38, 40, 43, 41], which explicitly takes the bias into account by incorporating its derivatives over the parameters of interest. Such approach has been successfully used for simple 1-parameter models in quantum imaging and metrology: estimation of the separation between 2 incoherent point sources [29, 37, 27] and inference of the wavelength by a spectrometer [24]. Minimization of the predicted estimation error over the bias allows to construct uniform (universal) version of CRB [45, 46, 47]. The resulting trade-off between the bias and the variance of the estimate can serve for resolution quantification in biomicroscopy [48, 49, 50]. While the approach is fruitful for derivation of general theoretical bounds on estimation errors, its practical application to quantifying information content of particular quantum imaging methods and setups can be cumbersome. Analytical calculation of neither bias gradient for a selected estimator, nor the optimal bias seems feasible for a real multiparametric problem of object reconstruction in quantum imaging [28], thus leading to the necessity of time-consuming numerical Monte-Carlo simulations.
@@ II.1 Fisher information
General purpose of measurements, including the ones used in imaging, is to estimate physical characteristics of the investigated system from the collected data [54]. For that purpose, one typically needs to construct the theoretical model of the signal $\mathbf{S}(\boldsymbol{\theta})$, where the vector $\boldsymbol{\theta}=\{\theta_{\mu}\}$ describes the parameters to be inferred from the experiment. Actually collected signal
@@
where the symbol $\operatorname{E}$ denotes mathematical expectation over the distribution of the random signal $\mathbf{Y}$. Roughly speaking, larger values of the FIM elements correspond to more informative measurements.
The significance of FIM is enhanced by its usage in Cramér-Rao bound (CRB) [21, 22]. If the parameters inference algorithm yields unbiased estimate $\hat{\boldsymbol{\theta}}$, i.e. its mathematical expectation equals the true value of the parameters: $\operatorname{E}(\hat{\boldsymbol{\theta}})=\boldsymbol{\theta}$, the covariance matrix of the estimated parameters is bounded from below by the inverse of the FIM:
@@
Eq. (3) can be also rewritten as the expectation value of the log-likelihood curvature (see e.g. [54]):
@@ II.2 Fisher information in quantum imaging
where $E^{(\pm)}(\mathbf{r},t)$ are positive(negative)-frequency field operators. Practically, the $n$-th order correlation function $G^{(n)}$ describes the rate of $n$-photon coincidence events for specified detection positions and time delays. If the main source of the detected signal fluctuations is the shot noise, the expression (3) can be simplified and takes the form (see, e.g. [33, 30, 66, 34]):
$F_{\mu\nu}=\sum_{i}\frac{1}{S_{i}(\boldsymbol{\theta})}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\mu}}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\nu}},$ (8)
where $S_{i}(\boldsymbol{\theta})$ are components of the signal $\mathbf{S}(\boldsymbol{\theta})$, represented as a vector $\mathbf{S}(\boldsymbol{\theta})=(S_{1}(\boldsymbol{\theta}),S_{2}(\boldsymbol{\theta}),\ldots)$.
@@
The close connection of FI to achievable resolution allows to use it for optimization of the imaging setup. Increase of FI, including removal of “Rayleigh’s curse” in simple cases, indicated apparent advantages of detection protocols based on spatial modes demultiplexing [29, 30, 33, 73, 74], modification of point-spread function (PSF) [31, 32], or probabilistic transformation of an entangled light state [34]. Analysis of FI proved to be useful in optimization of quantum light sources [28, 34] and selection of the optimal cumulant order in super-resolution optical fluctuation imaging (SOFI) [35, 36].
@@ III.2 Effective Fisher information matrix for constrained estimation
For unconstrained problems, FIM can be used for efficient prediction of experimental inference inaccuracies via CRB (Eq. (4)). In high-SNR regime, FIM also represents the quadratic form kernel for Gaussian approximation of the estimated parameters distribution (Eq. (6)). Mathematically, the two meanings of FIM are ensured by the relation between the quadratic form and the covariance matrix of a multivariate normal distribution:
@@
Figure 4: Scheme of taking constraints into account by construction of modified FIM. The basic FIM $F$ corresponds to the unconstrained problem and, in high-SNR regime, represents the quadratic form kernel for the probability distribution $p(\hat{\boldsymbol{\theta}})$ of unconstrained parameter estimates. Then, the constraints are applied to the probability distribution and the result is approximated by a multivariate normal distribution $\tilde{p}(\hat{\boldsymbol{\theta}})$. The kernel $\tilde{F}$ of that distribution’s quadratic form can be interpreted as the modified FIM since it is related to the covariance of the constrained parameter estimates through usual CRB.
@@
Then, the $(i+1)$-th iteration transforms $p^{(i)}(\hat{\boldsymbol{\theta}})$ into $p^{(i+1)}(\hat{\boldsymbol{\theta}})$ and includes the following steps [53] (see Appendix B for a visual illustration and additional expressions):
  • Construct a matrix $T=T^{T}=(F^{(i)})^{1/2}$ and apply a linear coordinate transformation $\hat{\boldsymbol{\theta}}\mapsto\boldsymbol{\theta}^{\prime}=T(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}^{(i)})$, such that the transformed distribution has zero mean ${\boldsymbol{\theta}^{(i)}}^{\prime}=0$ and the quadratic form kernel represented by the unit matrix: ${F^{(i)}}^{\prime}=T^{-1}F^{(i)}T^{-1}=1$. The transformation is non-degenerate, if $F^{(i)}$ is not singular.
@@
  • Shrink the distribution in the direction $\mathbf{d}=\mathbf{a}_{j}^{\prime}/|\mathbf{a}_{j}^{\prime}|$ corresponding to the found constraint: ${F^{(i)}}^{\prime}\mapsto{F^{(i)}}^{\prime\prime}+\xi\mathbf{d}\mathbf{d}^{T}$, ${\boldsymbol{\theta}^{(i)}}^{\prime}\mapsto{\boldsymbol{\theta}^{(i)}}^{\prime\prime}=-\delta\mathbf{d}$. Explicit expressions for $\delta$ and $\xi$ are listed in Appendix B.
@@ III.3 Ill-defined Fisher information matrix for dark objects
where $A(\mathbf{s})$ is the transmission amplitude of the object at the point with the transverse position $\mathbf{s}$, and $h(\mathbf{s},\mathbf{r})$ is the PSF (Green’s function) of the optical system. For a pixelized object and an array detector, the second order correlation between the $i$-th and $j$-th pixels of the detector can be expressed as [28]
$G^{(2)}_{ij}\propto|\Psi_{ij}|^{2}=\left|\sum_{m}D^{(ij)}_{m}A_{m}^{2}\right|^{2}$ (15)
with certain coefficients $D^{(ij)}_{m}$ (real-valued for simple models [28]). According to Eq. (8), diagonal elements of FIM are proportional to
$F_{mm}\propto A_{m}^{2}\sum_{i,j}\left(D^{(ij)}_{m}\right)^{2}\propto A_{m}^{2}.$ (16)
@@ III.4 Regularization of ill-defined Fisher information matrix
will be exactly the half-width of the peak (Fig. 5(b)). For the function $y_{2}(x)$, Eq. (19) provides an approximate result, which, however, is still quite accurate (see Fig. 5(c) and Appendix C).
@@
One should treat the discussed illustrative example with caution: the toy example deals with a continuous dependence $y(x)$, while the actually detected signal $\mathbf{Y}$ in quantum imaging typically represents a vector of discrete integer-valued random variables. Moreover, it is the discretness of the detected signal that leads to singularity of FIM for a dark object (Appendix D). Still, the gained insights can be used for regularization of FIM.
@@ IV.1 One-parameter model
The likelihood for the signal $Y$ (taking integer values) corresponds to Poisson distribution:
$L(Y|S(A))=\frac{S^{Y}(A)}{Y!}e^{S(A)}.$ (29)
Eq. (8) for FIM is applicable in the considered case and yields the result
@@
It is worth noting that the latter one also represents the optimal biased estimate (following the spirit of Refs. [44, 43, 47]), minimizing the mean squared error (MSE) averaged over the range of the true values $A\in[0,1]$ (Appendix E.1).
Fig. 6(a) shows the bias of the estimates $\hat{A}_{\text{MLE}}$ and $\hat{A}_{\text{B}}$ for $N=200$, $\eta=0.7$, $n=2$. One can see that the estimates are, indeed, biased near the boundary values $A=0$ and $A=1$ (in the regions I and III respectively).
@@
Fig. 6(b) shows the dependence of the estimation error
@@
on the true value $A$ of the transmission amplitude for the two estimators (blue and orange lines correspond to MLE and Bayesian estimators respectively). For each estimator, the MSE is calculated directly from the probability distribution (solid colored lines) and from modified CRB taking the bias into account [43] (dashed colored lines). One can notice that in the region I the error strongly depends on the estimator used; changing the bias, one can “redistribute” the inaccuracy over the range of the parameter $A$ [41]. Still, the error remains finite for the whole range $A\in[0,1]$. The details of the calculations are described in Appendix E.2.
Black lines in Fig. 6(b) correspond to application of basic CRB (namely, Eq. (20)) to the unmodified FI $F(A)$ (solid black line), FI after regularization according to Eq. (23) (dotted black line; see Appendix E.3), and FI $\tilde{F}(A)$ regularized and then corrected for the constraints (dot-dashed black line; see Appendix E.4). The regularization procedure resolves the divergence of the basic CRB with unmodified FI at $A=0$. The correction, described in Subsection III.2, further improves the error estimate and makes it consistent with the actual inaccuracies indicated by the colored lines in Fig. 6(b) (up to the strong variation of the error for different estimators in the region I).

### v3

@@ Abstract
Advanced super-resolution imaging techniques require specific approaches for accurate and consistent estimation of the achievable spatial resolution. Fisher information supplied to Cramér-Rao bound (CRB) has proved to be a powerful and efficient tool for resolution analysis and optical setups optimization. However, the standard CRB is not applicable to constrained problems violating the unbiasedness condition, while such models are frequently encountered in quantum imaging of complex objects. Complementary to the existing approaches based on modifying CRB, we propose a practical algorithm for approximate construction of a modified Fisher information matrix, which takes the constraints into account and enables accurate estimation of errors for constrained problems by the standard CRB. We demonstrate the efficiency of the proposed technique by applying it to 1-, 2-, and multi-parameter model problems in quantum imaging. The approach provides quantitative explanation of previous results with successful experimental reconstruction of objects with the spatial scale smaller than the theoretical limit predicted by the standard CRB.
@@ I Introduction
Efficiency of CRB for resolution analysis on the one hand and typicality of constrained problems in practical imaging on the other hand motivate one to search for a way to restore applicability of CRB to such problems. One of the available approaches is based on constructing the modified CRB [38, 40, 43, 41], which explicitly takes the bias into account by incorporating its derivatives over the parameters of interest. Such approach has been successfully used for simple 1-parameter models in quantum imaging and metrology: estimation of the separation between 2 incoherent point sources [29, 37, 27] and inference of the wavelength by a spectrometer [24]. Minimization of the predicted estimation error over the bias allows us to construct uniform (universal) version of CRB [45, 46, 47]. The resulting trade-off between the bias and the variance of the estimate can serve for resolution quantification in biomicroscopy [48, 49, 50]. While the approach is fruitful for derivation of general theoretical bounds on estimation errors, its practical application to quantifying information content of particular quantum imaging methods and setups can be cumbersome. Analytical calculation of neither bias gradient for a selected estimator, nor the optimal bias seems feasible for a real multiparametric problem of object reconstruction in quantum imaging [28], thus leading to the necessity of time-consuming numerical Monte-Carlo simulations.
@@ II.1 Fisher information
General purpose of measurements, including those used in imaging, is to estimate physical characteristics of the investigated system from the collected data [54]. For that purpose, one typically needs to construct the theoretical model of the signal $\mathbf{S}(\boldsymbol{\theta})$, where the vector $\boldsymbol{\theta}=\{\theta_{\mu}\}$ describes the parameters to be inferred from the experiment. Actually collected signal
@@
where the symbol $\operatorname{E}$ denotes mathematical expectation over the distribution of the random signal $\mathbf{Y}$. Roughly speaking, larger values of the FIM elements correspond to more informative measurements. If the signal $\mathbf{Y}$ takes discrete values, Eq. (3) can be re-written in the following way convenient for direct calculation:
$F_{\mu\nu}=\sum_{\mathbf{Y}}\frac{1}{L(\mathbf{Y}|\mathbf{S}(\boldsymbol{\theta}))}\frac{\partial L(\mathbf{Y}|\mathbf{S}(\boldsymbol{\theta}))}{\partial\theta_{\mu}}\frac{\partial L(\mathbf{Y}|\mathbf{S}(\boldsymbol{\theta}))}{\partial\theta_{\nu}},$ (4)
where the summation is performed over all possible realizations of the signal $\mathbf{Y}$.
The significance of FIM is enhanced by its usage in Cramér-Rao bound (CRB) [21, 22]. If the parameters inference algorithm yields unbiased estimate $\hat{\boldsymbol{\theta}}$, i.e., its mathematical expectation equals the true value of the parameters: $\operatorname{E}(\hat{\boldsymbol{\theta}})=\boldsymbol{\theta}$, the covariance matrix of the estimated parameters is bounded from below by the inverse of the FIM:
@@
Equation (3) can be also rewritten as the expectation value of the log-likelihood curvature (see e.g., Ref. [54]):
@@ II.2 Fisher information in quantum imaging
where $E^{(\pm)}(\mathbf{r},t)$ are positive(negative)-frequency field operators. Namely, the signal $\mathbf{S}(\boldsymbol{\theta})=(S_{1}(\boldsymbol{\theta}),S_{2}(\boldsymbol{\theta}),\ldots)$ is composed of the components
$S_{i}(\boldsymbol{\theta})\propto G^{(n)}(\mathbf{r}_{1}^{(i)},t_{1}^{(i)};\ldots;\mathbf{r}_{n}^{(i)},t_{n}^{(i)})$ (9)
defined by specific arguments of the correlation function.
From practical point of view, the $n$-th order correlation function $G^{(n)}$ describes the rate of $n$-photon coincidence events for specified detection positions and time delays. The signal component $S_{i}(\boldsymbol{\theta})$ describes the mean number of the corresponding coincidence events, while the component $Y_{i}$ of a random signal realization $\mathbf{Y}=(Y_{1},Y_{2},\ldots)$ equals the realized (integer) number of such events.
Commonly, the detection events can be treated as independent ones and the likelihood function corresponds to independent random variables $Y_{i}$ obeying Poisson distribution with the mean values $S_{i}(\boldsymbol{\theta})$:
$L(\mathbf{Y}|\mathbf{S}(\boldsymbol{\theta}))=\prod_{i}L(Y_{i}|S_{i}(\boldsymbol{\theta})),$ (10)
where
$L(Y|S(\boldsymbol{\theta}))=\frac{[S(\boldsymbol{\theta})]^{Y}}{Y!}e^{-S(\boldsymbol{\theta})}$ (11)
is the likelihood function for a single-variate Poisson distribution.
In that specific, but commonly encountered case the expression (4) can be simplified and takes the following form (see Refs. [33, 30, 66, 34] and Appendix B):
$F_{\mu\nu}=\sum_{i}\frac{1}{S_{i}(\boldsymbol{\theta})}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\mu}}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\nu}}.$ (12)
It is worth noting that the structure of Eqs. (4) and (12) for the FIM elements $F_{\mu\nu}$ is very similar, while their meaning is quite different. The general expression (4) is constructed from probabilities of all possible (mutually exclusive) measurement outcomes $\mathbf{Y}$, and the normalization $\sum_{\mathbf{Y}}L(\mathbf{Y}|\mathbf{S}(\boldsymbol{\theta}))=1$ holds. In Eq. (12), the summation index $i$ enumerates different signal components measured simultaneously in one experiment. The quantities $S_{i}(\boldsymbol{\theta})$ represent the expectation values of the events numbers and are not subjected to any specific kind of normalization.
@@
The close connection of FI to achievable resolution allows us to use it for optimization of the imaging setup. Increase of FI, including removal of “Rayleigh’s curse” in simple cases, indicated apparent advantages of detection protocols based on spatial modes demultiplexing [29, 30, 33, 73, 74], modification of point-spread function (PSF) [31, 32], or probabilistic transformation of an entangled light state [34]. Analysis of FI proved to be useful in optimization of quantum light sources [28, 34] and selection of the optimal cumulant order in super-resolution optical fluctuation imaging (SOFI) [35, 36].
@@ III.2 Effective Fisher information matrix for constrained estimation
For unconstrained problems, FIM can be used for efficient prediction of experimental inference inaccuracies via CRB (Eq. (5)). In high-SNR regime, FIM also represents the quadratic-form kernel for Gaussian approximation of the estimated parameters distribution (Eq. (7)). Mathematically, the two meanings of FIM are ensured by the relation between the quadratic form and the covariance matrix of a multivariate normal distribution:
@@
Figure 4: Scheme of taking constraints into account by construction of modified FIM. The basic FIM $F$ corresponds to the unconstrained problem and, in high-SNR regime, represents the quadratic-form kernel for the probability distribution $p(\hat{\boldsymbol{\theta}})$ of unconstrained parameter estimates. Then, the constraints are applied to the probability distribution and the result is approximated by a multivariate normal distribution $\tilde{p}(\hat{\boldsymbol{\theta}})$. The kernel $\tilde{F}$ of that distribution’s quadratic form can be interpreted as the modified FIM since it is related to the covariance of the constrained parameter estimates through usual CRB.
@@
Then, the $(i+1)$-th iteration transforms $p^{(i)}(\hat{\boldsymbol{\theta}})$ into $p^{(i+1)}(\hat{\boldsymbol{\theta}})$ and includes the following steps [53] (see Appendix C for a visual illustration and additional expressions):
  • Construct a matrix $T=T^{T}=(F^{(i)})^{1/2}$ and apply a linear coordinate transformation $\hat{\boldsymbol{\theta}}\mapsto\boldsymbol{\theta}^{\prime}=T(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}^{(i)})$, such that the transformed distribution has zero mean ${\boldsymbol{\theta}^{(i)}}^{\prime}=0$ and the quadratic-form kernel represented by the unit matrix: ${F^{(i)}}^{\prime}=T^{-1}F^{(i)}T^{-1}=1$. The transformation is non-degenerate, if $F^{(i)}$ is not singular.
@@
  • Shrink the distribution in the direction $\mathbf{d}=\mathbf{a}_{j}^{\prime}/|\mathbf{a}_{j}^{\prime}|$ corresponding to the found constraint: ${F^{(i)}}^{\prime}\mapsto{F^{(i)}}^{\prime\prime}+\xi\mathbf{d}\mathbf{d}^{T}$, ${\boldsymbol{\theta}^{(i)}}^{\prime}\mapsto{\boldsymbol{\theta}^{(i)}}^{\prime\prime}=-\delta\mathbf{d}$. Explicit expressions for $\delta$ and $\xi$ are listed in Appendix C.
@@ III.3 Ill-defined Fisher information matrix for dark objects
where $A(\mathbf{s})$ is the transmission amplitude of the object at the point with the transverse position $\mathbf{s}$, and $h(\mathbf{s},\mathbf{r})$ is the PSF (Green’s function) of the optical system.
For a pixelized object, an array detector, and a general case of non-ideally correlated biphotons, the second order correlation between the $i$-th and $j$-th pixels of the detector can be expressed as [28]
$G^{(2)}_{ij}\propto|\Psi_{ij}|^{2}=\biggl{|}\sum_{m,l}D^{(ij)}_{ml}A_{m}A_{l}\biggr{|}^{2}$ (19)
with certain coefficients $D^{(ij)}_{ml}$ (real-valued for simple models [28]). If biphotons are ideally correlated, the additional relation $D^{(ij)}_{ml}=0$ holds for $m\neq l$. According to Eq. (12), diagonal elements of FIM are proportional to
$F_{mm}\propto A_{m}^{2}\sum_{i,j}\left(D^{(ij)}_{mm}\right)^{2}\propto A_{m}^{2}.$ (20)
@@ III.4 Regularization of ill-defined Fisher information matrix
will be exactly the half-width of the peak (Fig. 5(b)). For the function $y_{2}(x)$, Eq. (23) provides an approximate result, which, however, is still quite accurate (see Fig. 5(c) and Appendix D).
@@
One should treat the discussed illustrative example with caution: the toy example deals with a continuous dependence $y(x)$, while the actually detected signal $\mathbf{Y}$ in quantum imaging typically represents a vector of discrete integer-valued random variables. Moreover, it is the discretness of the detected signal that leads to singularity of FIM for a dark object (Appendix E). Still, the gained insights can be used for regularization of FIM.
@@ IV.1 One-parameter model
The likelihood for the signal $Y$ (taking integer values) corresponds to Poisson distribution (11) with the parameters vector $\boldsymbol{\theta}$ containing the only component $A$. Equation (12) for FIM is applicable in the considered case and yields the result
@@
It is worth noting that the considered toy problem can be parameterized in different ways. For example, one can choose the probability of photon transmission $T=A^{2}$ as the parameter of interest instead of the transmission amplitude $A$ and get a different expression for the FI (Appendix F). However, such parameterization would be quite unnatural for a more general multi-parameter case with light interference effects: for example, the transmission amplitude directly enters the expression (19) for imaging with non-ideal spatial correlations of biphotons. The main goal of the current subsection is illustration of the developed approach by a simple example and preparation for analysis of practical multiparametric problems in Subsection IV.3. For that reason, we follow the general course here and use $A$ as the parameter of interest, rather than search for the optimal parameterization of the specific toy problem. Further discussion of parameterization influence on FI is provided in Appendix F.
@@
It is worth noting that the latter one also represents the optimal biased estimate (following the spirit of Refs. [44, 43, 47]), minimizing the mean squared error (MSE) averaged over the range of the true values $A\in[0,1]$ (Appendix G.1).
The bias of the estimates $\hat{A}_{\text{MLE}}$ and $\hat{A}_{\text{B}}$ is shown in Fig. 6(a) for $N=200$, $\eta=0.7$, $n=2$. One can see that the estimates are, indeed, biased near the boundary values $A=0$ and $A=1$ (in the regions I and III respectively).
@@
Figure 6(b) shows the dependence of the estimation error
@@
on the true value $A$ of the transmission amplitude for the two estimators (blue and orange lines correspond to MLE and Bayesian estimators respectively). For each estimator, the MSE is calculated directly from the probability distribution (solid colored lines) and from modified CRB taking the bias into account [43] (dashed colored lines). The details of the calculations are described in Appendix G.2.
One can notice that in the region I the error strongly depends on the estimator used; changing the bias, one can “redistribute” the inaccuracy over the range of the parameter $A$ [41]. MLE is heavily biased to zero: it maps zero signal $Y=0$ to $\hat{A}_{\text{MLE}}(0)=0$, while such outcome remains the most probable one for the whole range $A\in[0,A_{0}]$ with $A_{0}=N^{-1/(2n)}\eta^{-1/2}=0.32$. The estimation error is vanishing for $A=0$, but becomes quite large around the value $A=A_{0}$. On the opposite, the Bayesian estimate takes into account that zero signal is probable for $A\in[0,A_{0}]$ and maps it to $A_{\text{B}}(0)\approx A_{0}\Gamma(1/n)/\Gamma(1/(2n))=0.49A_{0}$. It is worth also mentioning the modified MLE, introduced in Appendix 5 of Ref. [30], where the MLE mapping is overridden by $A_{\text{MLE}}^{\prime}=A_{0}/2$ to improve the worst-case estimation error. Still, the MSE remains finite for the whole range $A\in[0,1]$ for the considered estimators.
Black lines in Fig. 6(b) correspond to application of basic CRB (namely, Eq. (24)) to the unmodified FI $F(A)$ (solid black line), FI after regularization according to Eq. (27) (dotted black line; see Appendix G.3), and FI $\tilde{F}(A)$ regularized and then corrected for the constraints (dot-dashed black line; see Appendix G.4). The regularization procedure resolves the divergence of the basic CRB with unmodified FI at $A=0$. The correction, described in Subsection III.2, further improves the error estimate and makes it consistent with the actual inaccuracies indicated by the colored lines in Fig. 6(b) (up to the strong variation of the error for different estimators in the region I).